Tohban Report 2008-12-03
From RHESSI Wiki
|Start Date:||26 Nov 2008|
|End Date:||3 Dec 2008|
|Next Tohban:||Jim McTiernan|
|List all reports|
No flares this week; GOES activity has been completely flat for the last few days.
The SSR fill is high, reaching 30-50% after each of the last few passes. This is likely a temporary problem due to having ~8 passes per day instead of the usual 10.
The spacecraft was spun up from 14.6 to 15 rpm, starting at 1:30 PST on Dec. 2nd.
At 3:38:00 UT today, decimation was begun for the northern magnetic zones. Decimation is no longer taking place in the southern magnetic zones.
The temperature of cold plate 2 is dropping, currently at 95.7K.
Monitor rate gaps:
- 2008-11-28T13:40:00.000 -- 2008-11-28T14:05:00.000 1500.0000
- 2008-12-02T03:15:00.000 -- 2008-12-02T07:20:00.000 14700.000
- 2008-12-02T08:50:00.000 -- 2008-12-03T00:00:00.000 54600.000
The RHESSI Ops meeting next Wednesday (December 10) will take place at 1 pm PST instead of the usual 2 pm to accommodate the Solar Cycle 24 meeting schedule. Most participants will be phoning in from the meeting; no conference room will be scheduled at SSL.